40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 20 00:19:23 2013, rana, Update, LSC, PRMI / DRMI investigations 6x
    Reply  Tue Aug 20 01:28:47 2013, Jenne, Update, LSC, REFL investigations 
       Reply  Tue Aug 20 11:41:30 2013, Koji, Update, LSC, REFL investigations 
          Reply  Tue Aug 20 11:52:20 2013, Jenne, Update, LSC, REFL investigations 
             Reply  Tue Aug 20 18:42:57 2013, Jenne, Update, LSC, REFL investigations 
Message ID: 9037     Entry time: Tue Aug 20 00:19:23 2013     Reply to this: 9038
Author: rana 
Type: Update 
Category: LSC 
Subject: PRMI / DRMI investigations 

While Jenne was plotting, I locked and aligned the MICH with AS55_Q. Then I aligned the PRM and locked PRMI using REFL55_I/Q with triggering on POP22, but no power normalization.

I used this to set the phase for REFL11 and REFL55 (driving PRM at 111.3 Hz and minimizing the Q response using the DTT Sine Response tool). I flipped the sign on REFL11 by 

The REFL11 gain is ~50x larger than REFL55; this is with the 15 dB whitening gain on REFL55 and none for REFL11. What's going on here? The attached PDF shows the two time series with the free swinging PRMI and both phases set to ~ +/- 2 deg. The REFL55 signals have been scaled up by 50x.

So then we went in and looked at the RF signals at the demod boards. To do this we disconnected the RFPD test cables and hooked the RF Mon outputs into the 50 Ohm inputs on a scope. The following PNG images show the scope traces. The REFL11 (yellow) traces are too big!! See how small the REFL55 (green) are. REFL11 is saturating - need to fix.

TEK00000.PNGTEK00001.PNG

TEK00002.PNGTEK00003.PNG

Attachment 1: REFL.pdf  54 kB  | Hide | Hide all
REFL.pdf
Attachment 6: REFL-2.pdf  68 kB  Uploaded Tue Aug 20 02:07:39 2013  | Hide | Hide all
REFL-2.pdf
ELOG V3.1.3-